a helium ballon has a volume of 2.40 L at 25.8 degrees celcius and a pressure of 1.17 atm at sea level. The ballon is released and floats upward. At a certain the atmospheric pressure is 0.810 atm and the temperature is 12.0 degrees celcius. Calculate the volume of the ballon
Search instead for a helium ballon has a volume of 2.40 L at 25.8 degrees celcius and a pressure of 1.17 atm at sea level. The ballon is released and floats upward. At a certain the atmospheric pressure is 0.810 atm and the tempurater is 12.0 degrees celcius. Calculate the volume of the ballon
1
Expert's answer
2021-05-31T02:08:39-0400

"V_1=2.40 \;L \\

T_1=25.8 +273.15=298.95 \;K \\

p_1 = 1.17 \;atm \\

p_2=0.810 \;atm \\

T_2 = 12.0+273.15=285.15 \;K \\

\frac{p_1V_1}{T_1}=\frac{p_2V_2}{T_2} \\

V_2 = \frac{p_1V_1T_2}{T_1p_2} \\

= \frac{1.17 \times 2.40 \times 285.15}{298.95 \times 0.810} \\

= 0.574 \;L"

Answer: 0.574 L